# TASTE: Text-Aligned Speech Tokenization and Embedding for Spoken Language Modeling

### v2

@@ Abstract
Recent efforts target spoken language models (SLMs) that not only listen but also speak for more natural human-LLM interaction. Joint speech-text modeling is a promising direction to achieve this. However, the effectiveness of recent speech tokens for joint modeling remains underexplored. To address this, we introduce Text-Aligned Speech Tokenization and Embedding (TASTE), a method that directly addresses the modality gap by aligning speech token with the corresponding text transcription during the tokenization stage. We propose a method that can achieve this through a attention-based aggregation mechanism and with speech reconstruction as the training objective. We conduct extensive experiments and show that TASTE can preserve essential paralinguistic information while dramatically reducing the token sequence length. With TASTE, we perform straightforward joint spoken language modeling by using Low-Rank Adaptation on the pre-trained text LLM. Experimental results show that TASTE-based SLMs perform comparable to previous work on SALMON and StoryCloze; while significantly outperform other pre-trained SLMs on speech continuation across subjective and objective evaluations. To our knowledge, TASTE is the first end-to-end approach that utilizes a reconstruction objective to automatically learn a text-aligned speech tokenization and embedding suitable for spoken language modeling. Our demo, code, and model are available at https://mtkresearch.github.io/TASTE-SpokenLM.github.io.
@@ 1 Introduction
Figure 1: The concept overview. Conventional methods extract speech tokens solely from speech, inducing length-mismatch problem when conducting joint speech-text modeling. By taking dual modalities as input, we generate speech tokenization that is aligned with text, facilitating straightforward and effective joint modeling.
Learning a speech representation suitable for universal speech-processing tasks has long been a significant challenge [31, 48, 42]. Unlike text—which can be encoded discretely [39, 18, 19]—speech is a continuous waveform carrying layered information (acoustic, semantic, prosodic, etc.). Recent neural self-supervised learning (SSL) methods move beyond filter-banks and MFCCs to encode raw audio into compact, high-dimensional embeddings that excel on discriminative tasks such as automatic speech recognition (ASR), emotion recognition, and speaker verification. Despite these gains, learning representations for generative speech tasks remains an open and more complex problem that has begun to attract focused attention [45, 17, 44, 28].
@@
Among the generative speech tasks, spoken language modeling (SLM) is an intriguing direction, aiming to create models that can not only listen but also speak. Typically, building an SLM requires two stages: first, deriving speech tokenizations; second, training a language model on the speech tokens. For the speech tokens, previous approaches either apply SSL-based representations following by discretization techniques [14, 21, 32, 11] or reuse units from neural codec models like EnCodec and SoundStream [5, 49, 20, 41]. Although autoregressive modeling with these speech tokens shows great potential in text-to-speech (TTS) [45, 47], previous SLMs that model only speech tokens [21, 32] have been shown to lack semantic fidelity [22].
@@
To bridge this gap, one promising direction is to utilize text—which is rich in semantic—during spoken language modeling. TWIST [11] shows that SLMs can benefit from initializing with text LLMs. More recent work often conducts joint speech-text modeling on tokens of both modalities to facilitate the semantic coherence on the generated speech [33, 6, 46, 9]. Yet integrating text and speech tokens introduces a length-mismatch challenge, as speech token sequences are usually longer than their text counterparts. Common remedies may include interleaving speech and text tokens [33] or inserting padding to synchronize sequence lengths between modalities [6, 46, 9]. However, these methods require either additional speech-text alignment or heuristic rules to enable joint modeling.
@@
In this work, we introduce Text-Aligned Speech Tokenization and Embedding (TASTE), a special type of speech tokenization tailored for speech-text joint spoken language modeling. By acknowledging that the length mismatch introduces additional complexity in joint modeling, we develop our speech token to be aligned with its corresponding text transcription tokens. To achieve this, we first obtain the textual transcription of a speech with the ASR model; then we derive the speech token based on the transcription through a specialized cross-attention mechanism for speech reconstruction. Note that the full process can be accomplished in an end-to-end manner, with no explicit speech-text alignment required. Unlike previous speech tokens that are developed under a fixed stride with fixed down-sampling rate, our speech token has dynamic frequency as it is text-aligned. Figure 1 shows an overall concept of TASTE, illustrating how our tokenization allows straightforward joint modeling.
@@
To evaluate the effectiveness of TASTE, we first conduct extensive experiments on speech reconstruction. Our results on LibriSpeech [34] show that TASTE not only resynthesizes speech in high quality, but also retains similarity to the original speech. TASTE achieves high-end reconstruction at an extremely low bit rate ($\sim$150 bps); while the other comparable methods are often more than thousands of bps. More intriguingly, we demonstrate that TASTE allows simple text-aligned speech editing. By exchanging the partial text-aligned speech tokens from two different utterances with the same content, we demonstrate that the paralinguistic information such as duration and tone can be exchanged precisely following the words being exchanged, resulting in natural edited speech.
@@
On the other hand, we demonstrate that TASTE successfully allows effective spoken language modeling. We perform straightforward joint modeling with TASTE under Low-Rank Adaptation [15]. We first perform speech continuation experiments with 3-second speech prompts given. The evaluation is three-fold. We use GPT-4o for evaluating the semantic aspect; UTMOS [38] for the acoustic aspect; and the human listening test for the general evaluation. Results show that our SLMs not only generates natural, meaningful speech continuations, but also outperforms the other 7B pre-trained SLMs across all the continuation evaluation aspects with 1.3B parameters. We also evaluate our SLMs on two benchmarks, SALMON [25] and StoryCloze [11] and our results show that our SLMs achieve comparable performance compared to the other speech-text joint modeling methods. Moreover, we show that our pretrained SLM can perform spoken question answering under few-shot scenario.
In summary, we derive TASTE, a text-aligned speech tokenization that allows effective joint speech-text spoken language modeling. By aligning the speech tokenization with its text counterpart during the tokenization stage, TASTE enables straightforward modeling. To our best knowledge, we are the first one to utilize the reconstruction objective to automatically derive a text-aligned speech tokenization and embedding that is suitable for joint speech-text spoken language modeling. Our demo is available at https://mtkresearch.github.io/TASTE-SpokenLM.github.io.
@@ 2 Method
We propose text-aligned speech tokenization and embedding (TASTE) to facilitate effective joint speech-text spoken language modeling. Here, we first introduce how we derive our tokenization—TASTE—in Section 2.1, and then discuss how we use TASTE for spoken language modeling (§ 2.2).
@@ 2.1.1 TASTE Speech Tokenizer
In TASTE, the speech tokenizer, denoted as $\mathrm{Tokenizer}(\cdot)$, is designed to generate the text-aligned speech tokenization and embedding with the speech-text pair $X=(\bm{u},\bm{v})$ taken as input, where $\bm{v}$ represents the textual transcription of the speech utterance $\bm{u}$, which can be easily obtained through an automatic speech recognition (ASR) system. Recent developments in robust and efficient ASR ([35, 10]) allow us to focus on discussing how to derive the text-aligned speech token effectively by assuming that $\bm{v}$ is of sufficient quality. The TASTE speech tokenizer is composed of three major components: an encoder, an aggregator, and a quantizer.
The encoder $\mathrm{Encoder}(\cdot)$ contains $L$ layers of Transformer ([43]) encoder blocks and is used to extract high-dimensional speech representation. We employ the pre-trained Whisper ASR encoder [35] as our speech encoder, and it is frozen during training. For an input speech utterance $\bm{u}$, the encoder produces a sequence of hidden states from each layer $[\bm{h}^{(1)},\bm{h}^{(2)},\ldots,\bm{h}^{(L)}]$. In our experiments, we retain the last hidden layer representation $\bm{h}^{(L)}$ and the shallow representation $\bm{h}^{(l)}$ from the first half of the hidden representations of the encoder for later usage, denoted as:
@@
To make the speech representation $\bm{z}$ text-aligned, we conduct a simple yet effective attention mechanism based on the three inputs. Consider that the original multi-head attention in [43] is denoted as $\mathrm{MultiHead}(Q,K,V)$, our first layer attention in the aggregator takes:
${Q}=\text{text transcription }\bm{v},\quad{K}=\text{encoder last hidden }\bm{h}^{(L)},\quad{V}=\text{encoder shallow hidden }\bm{h}^{(l)}.$
By doing so, the length of our first multi-head attention output should follow the text transcription $\bm{v}$. Note that the query of the following layers becomes the output from the previous layer. In addition, intuitions of using the encoder’s last hidden representation as keys, and the shallow hidden representation as values can be described as follows: 1) In Transformer-based ASR models, the last hidden states often encode rich speech-text alignment cues; sometimes the cross-attention weight matrices can even be exploited as soft word-alignment maps [35, 10]. 2) The shallow representation has been shown to support high-quality speech reconstruction even when the quantization is applied [7, 8]. Based on the above observations, we design our aggregator that can use the soft attention maps obtained from last encoder representations and the text transcriptions, to aggregate the shallow encoder representations that is beneficial for high-end speech reconstruction.
@@ 2.1.2 TASTE Speech Decoder
The unit decoder $\mathrm{UnitDecoder}(\cdot)$ is a Transformer-based decoder that takes the text token sequence $\bm{v}$ and the aligned speech embedding $\hat{\bm{z}}$ as input and predicts the speech unit $\bm{y}$ for reconstruction:
@@
Note that the additional speaker embedding is also taken as input to facilitate global speaker voice control in our spoken language models [16]. After we generating the speech unit $\bm{y}$, we use a unit-to-speech vocoder to further transform the unit into the reconstructed speech.
@@ 2.1.3 Training Objective
Similar to other reconstruction-based speech tokens [51, 24], we derive TASTE by training it for speech resynthesis. To achieve this, we extract the speech unit $\bm{y}^{\text{target}}$ with length $T^{\prime}$ from the original speech $u$ as the target unit for our speech tokenizer and speech decoder. Given the text transcription $\bm{v}$, the TASTE speech embedding $\hat{\bm{z}}$, and the unit from the original speech $\bm{y}^{\text{target}}$ as the target, the speech reconstruction through the tokenizer and the unit decoder parametrized by $\theta$ under the next prediction schema can be considered as minimizing the cross-entropy loss below:
@@
On the other hand, we employ the quantization loss as well to tokenize the continuous representation $\bm{z}$ extracted from the encoder-aggregator. Following prior works [5, 49], given that $\bm{z}^{(r)}$ is the $r$-th residual and $\hat{\bm{z}}^{(r)}$ indicates the $r$-th quantized residual, the the commitment loss is defined as:
@@
Note that to allow gradient to back-propagate from the unit decoder through the tokenizer, the straight-through estimation technique is applied towards the quantization process during traning.
@@ 2.2 TASTE for Spoken Language Modeling
Next, we describe how we conduct effective spoken language modeling with TASTE. Following previous work [11, 33], we perform pre-training on speech data. The text transcription of the speech data is also used for joint speech-text pre-training of our text-aligned spoken language model (TASLM). Since TASTE tokenization already aligns with the text token sequence, we can conduct a straightforward joint modeling, as illustrated in Figure 1. To demonstrate the robustness of TASTE, we perform two types of text-aligned spoken language modeling. First, we build $\rm{TASLM}_{\text{token}}$ over our text-aligned speech token $\bm{q}$, discussed in Section 2.2.1. Then, we show how we build $\rm{TASLM}_{\text{emb}}$ with our text-aligned speech embedding $\hat{\bm{z}}$, detailed in Section 2.2.2.

### v3

@@ Abstract
Recent efforts target spoken language models (SLMs) that not only listen but also speak for more natural human-LLM interaction. Joint text-speech modeling is a promising direction to achieve this. However, the effectiveness of recent speech tokens for joint modeling remains underexplored. To address this, we introduce Text-Aligned Speech Tokenization and Embedding (TASTE), a method that directly addresses the modality gap by aligning speech token with the corresponding text transcription during the tokenization stage. We propose a method that can achieve this through an attention-based aggregation mechanism and with speech reconstruction as the training objective. We have conducted extensive experiments to demonstrate that TASTE can preserve essential paralinguistic information while dramatically reducing the token sequence length. Moreover, TASTE enables straightforward joint spoken language modeling by using Low-Rank Adaptation on the pre-trained text LLM. Our experimental results show that joint modeling with TASTE outperforms other pre-trained SLMs in tasks such as speech continuation and likelihood-based next-speech selection, showcasing its effectiveness. To our best knowledge, TASTE is the first end-to-end approach that utilizes a reconstruction objective to learn a joint tokenization and embedding tailored for text-speech spoken language modeling. Our demo, code, and models are available at https://mtkresearch.github.io/TASTE-SpokenLM.github.io.
@@ 1 Introduction
Spoken language modeling (SLM) is an intriguing direction nowadays that aims at creating models that can not only listen but also speak (Lakhotia et al., 2021; Nguyen et al., 2023; Défossez et al., 2024; Fang et al., 2024; Arora et al., 2025). Typically, building an SLM requires two stages: first, deriving speech tokenizations; second, training a language model based on the speech tokens. For the speech tokens, previous approaches either apply self-supervised learning (SSL) representations following by discretization techniques (Baevski et al., 2020; Lakhotia et al., 2021; Nguyen et al., 2023; Hassid et al., 2023) or reuse units from neural codec models like EnCodec and SoundStream (Défossez et al., 2023; Zeghidour et al., 2021; Kumar et al., 2023; Siuzdak et al., 2024). Although autoregressive modeling with these speech tokens shows great potential in text-to-speech (TTS) (Wang et al., 2023; Xin et al., 2024a; Kim et al., 2024; Chen et al., 2024), previous SLMs that model only speech tokens (Lakhotia et al., 2021; Nguyen et al., 2023) have been shown to lack semantic fidelity (Lin et al., 2024).
@@
To bridge this gap, one promising direction is to leverage text—which is rich in semantics—during spoken language modeling. TWIST (Hassid et al., 2023) shows that SLMs can benefit from initializing with text LLMs. Building on this idea, recent work has shifted toward joint text–speech modeling to enhance semantic coherence in generated speech. Such approaches typically adopt a hybrid decoding scheme that generates both text and speech tokens. However, combining the two modalities introduces a length mismatch, since speech token sequences are usually much longer than their textual counterparts. Common remedies include interleaving text and speech tokens (Nguyen et al., 2025) or inserting padding to synchronize sequence lengths (Défossez et al., 2024; Xie and Wu, 2024a; Fang et al., 2024; Xie and Wu, 2024b), but these solutions rely on additional alignment procedures or heuristic rules, making joint modeling more complex.
As hybrid text–speech decoding becomes the prevailing paradigm for joint SLM (Défossez et al., 2024; Xie and Wu, 2024a; Fang et al., 2024; Li et al., 2025; Xie and Wu, 2024b), the design of speech tokens should be reconsidered in light of this setting. This motivates the development of more effective joint tokenization methods, which can be derived under the following two considerations: 1) a speech token should avoid redundantly encoding text content—already captured by the text tokens—and instead focus on conveying paralinguistic information; and 2) a straightforward one-to-one correspondence between text and speech tokens is preferable, as it allows SLMs to generate a text token and a speech token simultaneously without any heuristics or explicit alignments applied, mitigating the length mismatch issue during the tokenization stage.
Figure 1: The concept overview. Conventional methods extract speech tokens solely from speech, inevitably carries overlapped information with text tokens and induces length-mismatch problem when conducting joint text-speech modeling. By taking dual modalities as input, we generate speech tokenization that is aligned with text, facilitating straightforward and effective joint modeling.
@@
In this work, we introduce Text-Aligned Speech Tokenization and Embedding (TASTE), a special type of joint tokenization tailored for text-speech joint spoken language modeling. By acknowledging that the length mismatch introduces additional complexity in joint modeling, we develop our speech token to be aligned with its corresponding text transcription tokens. To achieve this, we first obtain the textual transcription of a speech with the ASR model; then we derive the speech token based on the transcription through a specialized cross-attention mechanism for speech reconstruction. Note that the full process can be accomplished in an end-to-end manner, with no explicit speech-text alignment required. Unlike previous speech tokens that are developed under a fixed stride with fixed down-sampling rate, our speech token has dynamic frequency as it is text-aligned. Figure 1 shows an overall concept of TASTE, illustrating how our joint tokenization allows effective joint modeling.
@@
To evaluate the effectiveness of TASTE, we first conduct extensive experiments on speech reconstruction. Our results on LibriSpeech (Panayotov et al., 2015) show that TASTE not only resynthesizes speech in high quality, but also retains similarity to the original speech. TASTE achieves high-end reconstruction at an extremely low bit rate ($\sim$150 bps); while the other comparable methods are often more than thousands of bps. We attribute the efficiency to the involvement of text tokens during encoding and decoding, and our speech tokens focus on carrying paralinguistic information, which is backed up by the demonstration that TASTE allows simple text-aligned speech editing. By exchanging the partial text-aligned speech tokens from two different utterances with the same content, we demonstrate that the paralinguistic information such as duration and tone can be exchanged precisely following the words being exchanged, resulting in natural edited speech.
@@
On the other hand, we demonstrate that TASTE successfully allows effective spoken language modeling. We perform straightforward joint modeling with TASTE under Low-Rank Adaptation (Hu et al., 2021). We first perform speech continuation experiments with 3-second speech prompts given. The evaluation is three-fold. We use GPT-4o for evaluating the semantic aspect; UTMOS (Saeki et al., 2022) for the acoustic aspect; and the human listening test for the general evaluation. Results show that our SLMs not only generates natural, meaningful speech continuations, but also outperforms the other 7B pre-trained SLMs across all the continuation evaluation aspects with 1.3B parameters. We also evaluate our SLMs on two benchmarks, SALMON (Maimon et al., 2024) and StoryCloze (Hassid et al., 2023) and our results show that our SLMs achieve comparable performance compared to the other text-speech joint modeling methods. Moreover, we show that our pretrained SLM can perform spoken question answering under the few-shot scenario.
In summary, we derive TASTE, a specialized tokenization approach for text–speech spoken language modeling. By aligning speech tokens with their text counterparts, TASTE provides a simple yet effective form of joint tokenization. Our results highlight joint tokenization as a key factor in joint modeling, offering a new perspective that may foster further research into more effective designs.
2 Related Work
Recent SLMs often require speech tokenization to conduct language modeling with the next prediction objective as the text LLMs. Unlike text, the speech signal is continuous and lengthy, making it difficult to derive proper speech tokenization for spoken language modeling. Common approaches may utilize self-supervised learned (SSL) speech models followed by quantization techniques to extract speech tokens (Baevski et al., 2020; Hsu et al., 2021; Lakhotia et al., 2021; Hassid et al., 2023; Nguyen et al., 2025). In addition, audio or speech codec models have also been used for tokenization in recent SLMs (Zeghidour et al., 2021; Défossez et al., 2023; 2024; Zhang et al., 2024). These models are designed for resynthesis, where the speech decoders are jointly learned with the encoders.
With speech tokenization, GSLM (Lakhotia et al., 2021) first demonstrates the possibility of building an SLM that can generate speech. TWIST (Hassid et al., 2023) further shows that the SLM can benefit from initialization with the text-only LLM. With regard to the huge success of text-only LLMs, recent work shifts the focus towards joint speech-text modeling (Défossez et al., 2024; Xie and Wu, 2024a). To facilitate joint modeling, Spirit LM (Nguyen et al., 2025) adopts an interleaving strategy; moshi (Défossez et al., 2024) trains its own tokenizer with a reduced token frequency. Moreover, delayed or sequential generation are introduced for joint modeling (Xie and Wu, 2024a).
Despite the increasing demand of joint speech-text modeling, we do not find any work discussing the effectiveness of current speech tokenization for it. However, we have observed that recent work is trying to mitigate the modality gap by reducing the token frequency (Défossez et al., 2024; Zeng et al., 2024), conducting an additional training stage for text-speech alignment (Xie and Wu, 2024a), or leveraging hidden representations of the ASR encoder (Du et al., 2024a). This motivates us to design a speech tokenization that is directly aligned with its text counterpart, tackling the mismatch issue during the tokenization stage.
In TASTE, we utilize a specialized mechanism based on attention to aggregate the encoder representations. We clarify that the text-speech cross-attention mechanism has also been used for fine-grained control of TTS. More specifically, Chen and Rudnicky (2022) propose content-style cross-attention to indicate their text-speech cross-attention mechanism that enables style transfer in TTS. Although both utilize a specialized text-speech cross-attention mechanism, the design choices and problem formulations are completely different. We attribute of our main novelty to inventing a text-aligned speech tokenization and embedding for joint spoken language modeling, and the text-speech cross-attention mechanism is considered and shown to be a clean and effective way of achieving it.
@@ 3 Method
We propose text-aligned speech tokenization and embedding (TASTE) to facilitate effective joint speech-text spoken language modeling. Here, we first introduce how we derive our joint tokenization in Section 3.1, and then discuss how we use TASTE for spoken language modeling (§ 3.2).
@@ 3.1.1 TASTE Speech Tokenizer
In TASTE, the speech tokenizer, denoted as $\mathrm{Tokenizer}(\cdot)$, is designed to generate the text-aligned speech tokenization and embedding with the speech-text pair $X=(\bm{u},\bm{v})$ taken as input, where $\bm{v}$ represents the textual transcription of the speech utterance $\bm{u}$, which can be easily obtained through an automatic speech recognition (ASR) system. Recent developments in robust and efficient ASR (Radford et al., 2023; Gandhi et al., 2023) allow us to focus on discussing how to derive the text-aligned speech token effectively by assuming that $\bm{v}$ is of sufficient quality. The TASTE speech tokenizer is composed of three major components: an encoder, an aggregator, and a quantizer.
The encoder $\mathrm{Encoder}(\cdot)$ contains $L$ layers of Transformer (Vaswani et al., 2017) encoder blocks and is used to extract high-dimensional speech representation. We employ the pre-trained Whisper ASR encoder (Radford et al., 2023) as our speech encoder, and it is frozen during training. For an input speech utterance $\bm{u}$, the encoder produces a sequence of hidden states from each layer $\bm{h}^{(1)},\bm{h}^{(2)},\ldots,\bm{h}^{(L)}]$. In our experiments, we retain the last hidden representation $\bm{h}^{(L)}$ and the shallow representation $\bm{h}^{(l)}$ from the first half of the hidden representations of the encoder:
@@
To make the speech representation $\bm{z}$ text-aligned, we conduct a simple yet effective attention mechanism based on the three inputs. Consider that the original multi-head attention in Vaswani et al. (2017) is denoted as $\mathrm{MultiHead}(Q,K,V)$, our first layer attention in the aggregator takes:
[…]
By doing so, the length of our first multi-head attention output should follow the text transcription $\bm{v}$. Note that the query of the following layers becomes the output from the previous layer. In addition, intuitions of using the encoder’s last hidden representation as keys, and the shallow hidden representation as values can be described as follows: 1) In Transformer-based ASR models, the last hidden states often encode rich speech-text alignment cues; sometimes the cross-attention weight matrices can even be exploited as soft word-alignment maps (Radford et al., 2023; Gandhi et al., 2023). 2) The shallow representation has been shown to support high-quality speech reconstruction even when the quantization is applied (Du et al., 2024a; b). Based on the above observations, we design our aggregator that can use the soft attention maps obtained from the last encoder representations and the text transcriptions, to aggregate the shallow encoder representations that have been shown to be beneficial for high-end speech reconstruction.
@@ 3.1.2 TASTE Speech Decoder
The unit decoder $\mathrm{UnitDecoder}(\cdot)$ is a Transformer-based decoder that takes the text token sequence $\bm{v}$ and the aligned speech embedding $\hat{\bm{z}}$ as condition and predicts the speech unit $\bm{y}$ for reconstruction:
@@
Note that the additional speaker embedding is also taken as input to facilitate global speaker voice control in our spoken language models (Ju et al., 2024). After we generating the speech unit $\bm{y}$, we use a unit-to-speech vocoder to further transform the unit into the reconstructed speech.
@@ 3.1.3 Training Objective
Similar to other reconstruction-based speech tokens (Zhang et al., 2024; Liu et al., 2025), we derive TASTE by training it for speech resynthesis. To achieve this, we extract the speech unit $\bm{y}^{\text{target}}$ with length $T^{\prime}$ from the original speech $u$ as the target unit for our speech tokenizer and speech decoder. Given the text transcription $\bm{v}$, the TASTE speech embedding $\hat{\bm{z}}$, and the unit from the original speech $\bm{y}^{\text{target}}$ as the target, the speech reconstruction through the tokenizer and the unit decoder parametrized by $\theta$ under the next prediction schema can be considered as minimizing the cross-entropy loss below:
@@
On the other hand, we employ the quantization loss as well to tokenize the continuous representation $\bm{z}$ extracted from the encoder-aggregator. Following prior works (Défossez et al., 2023; Zeghidour et al., 2021), given that $\bm{z}^{(r)}$ is the $r$-th residual and $\hat{\bm{z}}^{(r)}$ indicates the $r$-th quantized residual, the the commitment loss is defined as:
@@ 3.2 TASTE for Spoken Language Modeling
Next, we describe how we conduct effective spoken language modeling with TASTE. Following previous work (Hassid et al., 2023; Nguyen et al., 2025), we perform pre-training on speech data. The text transcription of the speech data is also used for joint speech-text pre-training of our text-aligned spoken language model (TASLM). Since TASTE tokenization already aligns with the text token sequence, we can conduct a straightforward joint modeling, as illustrated in Figure 1. To demonstrate the robustness of TASTE, we perform two types of text-aligned spoken language modeling. First, we build $\rm{TASLM}_{\text{token}}$ over our text-aligned speech token $\bm{q}$, discussed in Section 3.2.1. Then, we show how we build $\rm{TASLM}_{\text{emb}}$ with our text-aligned speech embedding $\hat{\bm{z}}$, detailed in Section 3.2.2.